OpenStudy (anonymous):

what is the midpoint of 2 and 4

OpenStudy (ivettef365):

you have 2 3 4 what is between ?

HanAkoSolo (jamierox4ev3r):

@zkrup stop providing answers without a process!!!

OpenStudy (anonymous):

3

OpenStudy (anonymous):

@Jamierox4ev3r Oh crap I'm sorry Didn't realize there was actually a process to this. To simple for me.

HanAkoSolo (jamierox4ev3r):

provide a process first!

HanAkoSolo (jamierox4ev3r):

i guess you are right... but help them visualize, like draw a number line ofr something... whatever makes sense.

HanAkoSolo (jamierox4ev3r):

the goal is for the user to learn the process so she will end up learning. We provide a learning process

OpenStudy (anonymous):

probably 3

OpenStudy (anonymous):

M=a+b=2+4 2 2

OpenStudy (anonymous):

M=(1,2) right

OpenStudy (anonymous):

Help me

OpenStudy (ivettef365):

2+4 6 ---- = --- = ??? 2 2

OpenStudy (austinl):

Step 1: Determine the two numbers whose midpoint you want to find. For example, assume you want to find the midpoint between the numbers 2 and 32. Step 2: Add the two numbers together. Continuing the same example, 32 + 2 = 34. Step 3: Divide the product by 2. Continuing the same example, 34 / 2 = 17. This figure represents the midpoint between the two numbers.

OpenStudy (austinl):

Therefore, \(\Large{\frac{2+4}{2}=\frac{6}{2}=3}\)

OpenStudy (anonymous):

so my answers is 3

OpenStudy (anonymous):

so it will be the same for this one 2 and -5 right???

OpenStudy (anonymous):

and that answers is -1.5 right

OpenStudy (austinl):

\(\Large{\frac{2+(-5)}{2}}=-\frac{3}{2}\)

OpenStudy (anonymous):

so the answers is -1.5 right

OpenStudy (austinl):

I would tend to leave answers in exact form for mathematics courses, it is a good habit to get into. But yes... -1.5

OpenStudy (anonymous):

ok thank you

OpenStudy (austinl):

No problem.

OpenStudy (anonymous):

one more thing this question find the coordinates of the midpoint of H(0,0),X(8,4) will it be the same thing as last time